Accelerator Update for December 7, 2007 - December 10, 2007
Friday December 7
The day shift began with Operations monitoring the Recycler stash (203.1E10) and the Antiproton Source (Pbar) stack (12.6E10), with the Main Injector (MI), Booster, and Linac operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting on the next store, with NuMI in access, and with SciBooNE / MiniBooNE experiments taking beam.
Operations prepared the Meson area enclosure M01 for access at 8:43 AM.
Operations put the TeV through a dry squeeze at 9:05 AM.
Operations put the TeV into shot setup at 9:45 AM.
The Neutron Therapy Facility (NTF) reported at 11:05 AM that they had completed treatment on their last patient for the day.
Operations prepared Linac for access at 11:44 AM.
Operations put Pbar into standby at 11:47 AM.
Operations turned off MI and prepared the area for access at 12:56 PM.
At 2:14 PM, the TeV sector F4 LCW pump tripped off and caused a Tevatron Electron Lens (TEL1) to trip off. Operations contacted experts. CDF and D0 turned off their detectors until the pump can be turned on and TEL1 brought back on line.
A TEL expert turned TEL1 on at 3:04 PM.
D0 brought their detector back online at 3:06 PM.
Operations began turning MI on at 4:08 PM.
NuMI came out of access at 4:43 PM.
At 6:37 PM, TEL1 tripped off due to a vacuum pump trip. Everything reset.
Linac came out of access at 11:09 PM.
Saturday December 8
The midnight shift began with Operations monitoring store 5787 (48.29E30), the stash (0.0E10), and the stack (60.8E10), and with NuMI and Sci /MiniBooNE off due to Linac vacuum recovery.
At 12:05 AM, a Linac vacuum pump (IPBD2) tripped off. It reset, but the vacuum was bad. Linac technicians turned off the pump to let it cool down.
A Linac expert reported at 12:23 AM that he was going to run some study pulses of beam to the Momentum dump to see its affect on vacuum.
The Linac expert reported at 12:33 AM that there was no adverse effect on vacuum. Operations began sending beam to Pbar.
Operations sent beam to NuMI at 12:45 AM.
At 1:06 AM, KRF2 tripped off. It reset.
A Linac expert reported at 2:48 AM that for the time being they could supply beam to Pbar and NuMI, but not to Sci/MiniBooNE until the vacuum gets better. IPBD2 tripped off and was turned back on several times throughout the midnight shift.
Operations began sending beam to Sci/MiniBooNE at a reduced rate at 3:57 AM.
Operations reported at 7:57 AM that the Recycler had stashed antiprotons four times during the midnight shift.
At 10:20 AM, CDF reported that they had a power supply fail for their trigger system. This stopped them from taking data for about 75 minutes.
At 11:10 AM, a Linac RF station (LRF4) tripped off and would not reset. Operations called in an expert.
At 11:51 AM, a Linac expert reported that he was worried about the Linac vacuum and asked to halt beam to Sci/MiniBooNE.
At 12:28 PM, a TeV sector C49V separator sparked.
A Linac expert brought LRF4 back on line at 3:11 PM.
Operations prepared Linac for an access at 4:46 PM. Experts were in and out in about 15 minutes.
Operations resumed sending beam to Pbar, NuMI, and Sci/MiniBooNE at 5:38 PM.
Operations terminated store 5787 at 6 PM.
Operations put the TeV into shot setup at 6:24 PM.
At 7:56 PM, LRF4 tripped off. It reset.
Operations reported at 8:01 PM that the Recycler had stashed once in the last 12 hours.
Operations established store 5788 at 8:11 PM with an initial luminosity of 172.45E30.
At 8:54 PM, LRF4 tripped off. It reset.
A Booster expert began tuning at 9:12 PM.
At 11:49 PM, KRF4 tripped off and would not reset. Operations called in an expert.
Sunday December 9
The midnight shift began with Operations monitoring store 5788, a stash, and stack, and with NuMI and Sci/MiniBooNE off due to a problem with KRF4.
Linac resumed sending beam at 3:18 AM.
Linac experts brought KRF4 back online at 3:31 AM. They found a large feed forward signal, which could only be fixed by powering down the low level RF crate for a couple of minutes and then bringing it back up.
At 3:41 AM, KRF6 tripped off and would not reset. A Linac expert was still here. He had to replace the station’s shunt power supply.
Linac resumed sending beam at 5:03 AM.
Operations reported at 7:58 AM that the Recycler had stashed four times during the last 12 hours.
At 8:12 AM, LRF4 tripped off. It reset.
At 10:59 AM, KRF6 tripped off. It reset.
At 11:36 AM, operators inspected the TeV sector C2 wet engine. They reported that its bearings were bad.
At 11:44 AM, KRF7 tripped off. It reset.
Operations terminated store 5788 at 12:16 PM.
Operations prepared D0 for an access at 12:30 PM.
At 12:47 PM, KRF7 tripped off. It reset.
At 1:08 PM, KRF2 tripped off. Operators managed to reset the station by 1:39 PM.
Cryo system technicians began repairs on the C2 wet engine at 1:46 PM.
At 2:12 PM, LRF3 tripped off. It reset.
D0 came out of access at 3:08 PM.
The Duty Electrician reported at 5:24 PM that the reason why the sector F4 LCW pump was tripping off was due to a bad connection. He reterminated the connections. The pump was back on by 6:08 PM.
At 6:51 PM, LRF4 tripped off. It reset.
Operations reported at 8:01 PM that the Recycler had stashed twice during the last 12 hours.
At 8:38 PM, the Linac beam valves closed due to a vacuum problem. This cleared up and beam resumed about 15 minutes later.
At 9:01 PM, the Pelletron tripped off. Operations contacted an expert. Operators reset it.
At 9:04 PM, KRF2 tripped off. It reset.
Operations reported that the C2 wet engine repairs were done by 9:06 PM. The cryo system resumed normal temperatures about 30 minutes later.
Operations put the TeV through a dry squeeze at 10:31 PM.
Operations put the TeV into shot setup at 10:55 PM.
At 11:22 PM, a TeV sector D2 feeddown sextupole power supply tripped off on a fuse fault. This put shot setup on hold.
Monday December 10
The midnight shift began with operators working on the D2 sextupole power supply, with Operations monitoring a stash (75.7E10), with Pbar stacking (16.2E10), and with NuMI and Sci/MiniBooNE taking beam.
Operators reported at 12:09 AM that they had replaced the D2 sextupole power supply.
At 1:12 AM, Operations reported that the solenoid collector power supply for TEL2 was off and would not reset. Operators talked with experts and they said that this should not be a problem. TEL1 was on. One expert suggested Operations restart shot setup.
At 2:11 AM, LRF4 tripped off. It reset.
Operations put the TeV through a dry squeeze at 2:35 AM.
At 4:36 AM, KRF2 tripped off. It reset.
Operations turned off the NuMI beam at 5 AM in preparation for a scheduled access.
Operations established store 5791 at 5:02 AM with an initial luminosity of approximately 176.85E30.
At 6:55 AM, Operations reported that a TeV sector E48 kicker tripped off. Operators contacted an expert.
A TeV kicker expert brought the E48 kicker back online at 7:08 AM.
Switchyard (SY) experts began taking beam to the SY absorber at 7:45 AM.
Operations reported at 7:48 AM that the Recycler had stashed once during the last 12 hours.
The Plans for Monday
Linac experts have determined that the Momentum dump does have a vacuum leak. The plan for today is to stash, stack, and store. The Run Coordinator says that Linac will continue to run with this configuration until parts are fabricated to cap off the Momentum dump.
Comments and Suggestions
|last modified 12/10/2007 email Fermilab|
|Security, Privacy, Legal|